# MARS: Unleashing the Power of Variance Reduction for Training Large Models

### v2

@@ Appendix C Proof of Auxiliary Lemmas
$\mathbf{u}_{t}$ $=a_{1}\mathbf{u}_{t-1}+a_{2}\mathbf{m}_{t},$ (C.1)
$\mathbf{m}_{t}$ $=b_{1}\mathbf{u}_{t}+b_{2}\mathbf{m}_{t}.$ (C.2)
@@
$\mathbf{m}_{t}=a_{1}\mathbf{m}_{t-1}+(b_{1}a_{2}-a_{1}b_{2}+b_{2})\mathbf{m}_{t}+a_{1}b_{2}(\mathbf{m}_{t}-\mathbf{g}_{t-1}).$
Proof of Lemma C.1.
Substituting (C.1) into (C.2), we obtain
$\mathbf{m}_{t}=b_{1}\big{(}a_{1}\mathbf{u}_{t-1}+a_{2}\mathbf{m}_{t}\big{)}+b_{2}\mathbf{m}_{t}=a_{1}b_{1}\mathbf{u}_{t-1}+(b_{1}a_{2}+b_{2})\mathbf{m}_{t}.$ (C.3)
On the other hand, shifting the index of (C.2) by $1$, it holds that
$\mathbf{m}_{t-1}=b_{1}\mathbf{u}_{t-1}+b_{2}\mathbf{g}_{t-1}.$ (C.4)
Combining (C.3) and (C.4), we obtain the iterative update of $\mathbf{m}_{t}$ from its previous value $\mathbf{m}_{t-1}$ as:
$\mathbf{m}_{t}$ $=a_{1}\mathbf{m}_{t-1}-a_{1}b_{2}\mathbf{g}_{t-1}+(b_{1}a_{2}+b_{2})\mathbf{m}_{t}$
$=a_{1}\mathbf{m}_{t-1}+(b_{1}a_{2}-a_{1}b_{2}+b_{2})\mathbf{m}_{t}+a_{1}b_{2}(\mathbf{m}_{t}-\mathbf{g}_{t-1}).$
@@
C.2 Lemma C.2 and Proof
Lemma C.2.
@@
$\langle\mathbf{m}_{t},\mathbf{x}_{t}-\mathbf{x}_{t+1}\rangle$ $\geq\frac{\rho(1-\eta_{t}\lambda)}{\eta_{t}}\|\mathbf{x}_{t}-\mathbf{x}_{t+1}\|_{2}^{2}+\frac{\lambda}{2}\big{[}\mathbf{x}_{t}^{\top}\mathbf{H}_{t+1}\mathbf{x}_{t}-\mathbf{x}_{t+1}^{\top}\mathbf{H}_{t}\mathbf{x}_{t+1}-\sqrt{2(1-\beta_{2,t})}D^{2}\big{]}.$ (C.5)
Proof of Lemma C.2.
@@
$\geq\frac{\rho(1-\eta_{t}\lambda)}{\eta_{t}}\|\mathbf{x}_{t}-\mathbf{x}_{t+1}\|_{2}^{2}-\lambda\left\langle\mathbf{H}_{t}\mathbf{x}_{t+1},\mathbf{x}_{t}-\mathbf{x}_{t+1}\right\rangle,$ (C.6)
where the inequality follows from Assumption 4.3. By convexity of $h_{1}(\mathbf{x}):=\frac{1}{2}\mathbf{x}^{\top}\mathbf{H}_{t}\mathbf{x}$, we obtain
@@
We recall that $\mathbf{H}_{t}=\text{diag}(\sqrt{\widehat{\mathbf{v}}_{t}}+\epsilon)$ in Algorithm 2. Combining this with Lemma B.1, we derive
@@
C.3 Proof of Lemma B.1
Proof of Lemma B.1.
@@
$\|\mathbf{v}_{t}\|_{2}$ $=\Big{\|}\sum_{k=1}^{t}(1-\beta_{2,k})\tilde{\mathbf{c}}_{k}^{2}\prod_{j=k+1}^{t}\beta_{2,j}+\prod_{j=1}^{t}\beta_{2,j}\mathbf{v}_{0}\Big{\|}_{2}\leq\sum_{k=1}^{t}(1-\beta_{2,k})\prod_{j=k+1}^{t}\beta_{2,j}=\Big{(}1-\prod_{k=1}^{t}\beta_{2,k}\Big{)}\leq 1,$ (C.7)
@@
$\|\mathbf{m}_{t}\|_{2}$ $=\Big{\|}\sum_{k=1}^{t}(1-\beta_{1,k})\tilde{\mathbf{c}}_{k}\prod_{j=k+1}^{t}\beta_{1,j}+\prod_{j=1}^{t}\beta_{1,j}\mathbf{m}_{0}\Big{\|}_{2}\leq\sum_{k=1}^{t}(1-\beta_{1,k})\prod_{j=k+1}^{t}\beta_{1,j}=\Big{(}1-\prod_{k=1}^{t}\beta_{1,k}\Big{)}\leq 1.$ (C.8)
@@
C.4 Proof of Lemma B.2
Proof of Lemma B.2.
@@
$\mathbf{m}_{t+1}-\nabla F(\mathbf{x}_{t+1})$ $=(1-\beta_{1,t+1})\bigg{(}\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\bigg{)}+\beta_{1,t+1}\big{(}\mathbf{m}_{t}-\nabla F(\mathbf{x}_{t})$
@@
$\bm{\varepsilon}_{t+1}=(1-\beta_{1,t+1})\big{(}\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\big{)}+\beta_{1,t+1}\bm{\varepsilon}_{t}+\beta_{1,t+1}\big{(}\gamma_{t+1}\bm{\Delta}_{t}-\mathbb{E}\bm{\Delta}_{t}\big{)},$ (C.9)
where the expectation in the last term is taken over the randomness in $\bm{\xi}_{t+1}$. Taking squared norm over both sides of (C.9) and then take expectation over $\bm{\xi}_{t+1}$, we have
$\mathbb{E}\|\bm{\varepsilon}_{t+1}\|_{2}^{2}$ (C.10)
$=\mathbb{E}\|(1-\beta_{1,t+1})\big{(}\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\big{)}+\beta_{1,t+1}\bm{\varepsilon}_{t}+\beta_{1,t+1}\big{(}\gamma_{t+1}\bm{\Delta}_{t}-\mathbb{E}\bm{\Delta}_{t}\big{)}\|_{2}^{2}$
@@
$\qquad+\underbrace{\beta_{1,t+1}^{2}(\gamma_{t+1}-1)^{2}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}}_{\mbox{$II$}}$
$\qquad-2(1-\gamma_{t+1})\beta_{1,t+1}\underbrace{\mathbb{E}\bigg{\langle}\bm{\Delta}_{t},(1-\beta_{1,t+1})\big{(}\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\big{)}+\beta_{1,t+1}\bm{\varepsilon}_{t}+\beta_{1,t+1}\big{(}\bm{\Delta}_{t}-\mathbb{E}\bm{\Delta}_{t}\big{)}\bigg{\rangle}}_{\mbox{$III$}}.$ (C.11)
Using $G_{t}$ to represent
$G_{t+1}$ $:=(1-\beta_{1,t+1})\mathbb{E}\bigg{\langle}\bm{\Delta}_{t},\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\bigg{\rangle}+\beta_{1,t+1}\mathbb{E}\bigg{\langle}\bm{\Delta}_{t},\bm{\varepsilon}_{t}\bigg{\rangle},$
The above can be further summarized as
$\mathbb{E}\|\bm{\varepsilon}_{t+1}\|_{2}^{2}$ $=I+\underbrace{\beta_{1,t+1}^{2}(\gamma_{t+1}-1)^{2}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}}_{\mbox{$II$}}-2(1-\gamma_{t+1})\beta_{1,t+1}\underbrace{\Big{(}G_{t+1}+\beta_{1,t+1}\big{(}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}-\|\mathbb{E}\bm{\Delta}_{t}\|_{2}^{2}\big{)}\Big{)}}_{\mbox{$III$}}$
$=I+\beta_{1,t+1}^{2}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}\bigg{(}(1-\gamma_{t+1})^{2}-2(1-\gamma_{t+1})\frac{G_{t+1}+\beta_{1,t+1}\big{(}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}-\|\mathbb{E}\bm{\Delta}_{t}\|_{2}^{2}\big{)}}{\beta_{1,t+1}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}}\bigg{)}.$ (C.12)
Therefore, defining
$A_{t+1}:=\frac{G_{t+1}+\beta_{1,t+1}\big{(}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}-\|\mathbb{E}\bm{\Delta}_{t}\|_{2}^{2}\big{)}}{\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}}.$
We further have
$\mathbb{E}\|\bm{\varepsilon}_{t+1}\|_{2}^{2}$ $=I+\beta_{1,t+1}^{2}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}\bigg{(}(1-\gamma_{t+1})^{2}-2(1-\gamma_{t+1})\frac{A_{t+1}}{\beta_{1,t+1}}\bigg{)}$
$=I+\beta_{1,t+1}^{2}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}\bigg{(}1-\gamma_{t+1}-\frac{A_{t+1}}{\beta_{1,t+1}}\bigg{)}^{2}-\beta_{1,t+1}^{2}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}\bigg{(}\frac{A_{t+1}}{\beta_{1,t+1}}\bigg{)}^{2}$
$=I+\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}\big{(}\beta_{1,t+1}(1-\gamma_{t+1})-A_{t+1}\big{)}^{2}-\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}A_{t+1}^{2}$
@@
$\leq 2(1-\beta_{1,t+1})^{2}\mathbb{E}\|\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\|_{2}^{2}+2\beta_{1,t+1}^{2}\mathbb{E}\|\bm{\Delta}_{t}-\mathbb{E}\bm{\Delta}_{t}\|_{2}^{2}+\beta_{1,t+1}^{2}\mathbb{E}\|\bm{\varepsilon}_{t}\|_{2}^{2}.$ (C.13)
Utilizing the fact that $\mathbb{E}\|\bm{\Delta}_{t}-\mathbb{E}\bm{\Delta}_{t}\|_{2}^{2}\leq\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}$ and the $L$-smoothness of $f$, we conclude that
@@
$\leq 2(1-\beta_{1,t+1})^{2}\sigma^{2}+2\beta_{1,t+1}^{2}L^{2}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}+\beta_{1,t+1}^{2}\mathbb{E}\|\bm{\varepsilon}_{t}\|_{2}^{2}-M_{t+1}$
where
$M_{t+1}:=\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}A_{t+1}^{2}-\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}\Big{(}\beta_{1,t+1}(1-\gamma_{t+1})-A_{t+1}\Big{)}^{2}.$
Rearranging the terms finishes the proof. ∎
C.5 Proof of Lemma B.3
Proof of Lemma B.3.
Given the L-smoothness of $F(\mathbf{x})$ in Assuption 4.2, we have
@@
$=F(\mathbf{x}_{t})+\left\langle\mathbf{m}_{t},\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\rangle+\left\langle\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t},\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{L}{2}\cdot\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}.$ (C.14)
By definition of $\mathbf{m}_{t}$, Assumption 4.3 and the update rule of $\mathbf{x}_{t+1}$ in Algorithm 1, using Lemma 1 in Ghadimi and Lan (2013), we have
$\langle\mathbf{m}_{t},\mathbf{x}_{t}-\mathbf{x}_{t+1}\rangle$ $\geq\frac{\rho}{\eta_{t}}\|\mathbf{x}_{t}-\mathbf{x}_{t+1}\|_{2}^{2}.$ (C.15)
Bringing (C.15) into (C.14), we obtain
@@
The second inequality follows from applying both the Cauchy-Schwarz and Young’s inequalities. Moreover, the final inequality results from selecting $\eta_{t}$ to satisfy $\eta_{t}<\frac{\rho}{2L}$ This completes our proof. ∎
C.6 Proof of Lemma B.4
Proof of Lemma B.4.
Bringing (C.5) in Lemma C.2 into (C.14), we have
@@
$\qquad-\frac{\rho(1-\eta_{t}\lambda)}{\eta_{t}}\|\mathbf{x}_{t}-\mathbf{x}_{t+1}\|_{2}^{2}-\frac{\lambda}{2}\big{[}\mathbf{x}_{t}^{\top}\mathbf{H}_{t+1}\mathbf{x}_{t}-\mathbf{x}_{t+1}^{\top}\mathbf{H}_{t}\mathbf{x}_{t+1}-\sqrt{2(1-\beta_{2,t})}D^{2}\big{]}.$
@@
$\qquad-\frac{\lambda}{2}\big{[}\mathbf{x}_{t}^{\top}\mathbf{H}_{t+1}\mathbf{x}_{t}-\mathbf{x}_{t+1}^{\top}\mathbf{H}_{t}\mathbf{x}_{t+1}-\sqrt{2(1-\beta_{2,t})}D^{2}\big{]}.$
@@
$\qquad-\frac{\lambda}{2}\big{[}\mathbf{x}_{t}^{\top}\mathbf{H}_{t+1}\mathbf{x}_{t}-\mathbf{x}_{t+1}^{\top}\mathbf{H}_{t}\mathbf{x}_{t+1}-\sqrt{2(1-\beta_{2,t})}D^{2}\big{]}.$
$=F(\mathbf{x}_{t})+\frac{\eta_{t}}{\rho}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}-\frac{\rho}{4\eta_{t}}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}-\frac{\lambda}{2}\big{[}\mathbf{x}_{t}^{\top}\mathbf{H}_{t+1}\mathbf{x}_{t}-\mathbf{x}_{t+1}^{\top}\mathbf{H}_{t}\mathbf{x}_{t+1}\big{]}$
$\qquad+\frac{\lambda}{2}\sqrt{2(1-\beta_{2,t})}D^{2}.$
@@
C.7 Proof of Lemma B.5
Proof of Lemma B.5.

### v3

@@ Appendix D Proof of Auxiliary Lemmas
$\mathbf{u}_{t}$ $=a_{1}\mathbf{u}_{t-1}+a_{2}\mathbf{g}_{t},$ (D.1)
$\mathbf{m}_{t}$ $=b_{1}\mathbf{u}_{t}+b_{2}\mathbf{g}_{t}.$ (D.2)
@@
$\mathbf{m}_{t}=a_{1}\mathbf{m}_{t-1}+(b_{1}a_{2}-a_{1}b_{2}+b_{2})\mathbf{g}_{t}+a_{1}b_{2}(\mathbf{g}_{t}-\mathbf{g}_{t-1}).$
Proof of Lemma D.1.
Substituting (D.1) into (D.2), we obtain
$\mathbf{m}_{t}=b_{1}\big{(}a_{1}\mathbf{u}_{t-1}+a_{2}\mathbf{g}_{t}\big{)}+b_{2}\mathbf{g}_{t}=a_{1}b_{1}\mathbf{u}_{t-1}+(b_{1}a_{2}+b_{2})\mathbf{g}_{t}.$ (D.3)
On the other hand, shifting the index of (D.2) by $1$, it holds that
$\mathbf{m}_{t-1}=b_{1}\mathbf{u}_{t-1}+b_{2}\mathbf{g}_{t-1}.$ (D.4)
Combining (D.3) and (D.4), we obtain the iterative update of $\mathbf{m}_{t}$ from its previous value $\mathbf{m}_{t-1}$ as:
$\mathbf{m}_{t}$ $=a_{1}\mathbf{m}_{t-1}-a_{1}b_{2}\mathbf{g}_{t-1}+(b_{1}a_{2}+b_{2})\mathbf{g}_{t}$
$=a_{1}\mathbf{m}_{t-1}+(b_{1}a_{2}-a_{1}b_{2}+b_{2})\mathbf{g}_{t}+a_{1}b_{2}(\mathbf{g}_{t}-\mathbf{g}_{t-1}).$
@@
D.2 Lemma D.2 and Proof
Lemma D.2.
@@
$\langle\mathbf{m}_{t},\mathbf{x}_{t}-\mathbf{x}_{t+1}\rangle$ $\geq\frac{\rho(1-\eta_{t}\lambda)}{\eta_{t}}\|\mathbf{x}_{t}-\mathbf{x}_{t+1}\|_{2}^{2}+\frac{\lambda}{2}\big{[}\mathbf{x}_{t}^{\top}\mathbf{H}_{t+1}\mathbf{x}_{t}-\mathbf{x}_{t+1}^{\top}\mathbf{H}_{t}\mathbf{x}_{t+1}-\sqrt{2(1-\beta_{2,t})}D^{2}\big{]}.$ (D.5)
Proof of Lemma D.2.
@@
$\geq\frac{\rho(1-\eta_{t}\lambda)}{\eta_{t}}\|\mathbf{x}_{t}-\mathbf{x}_{t+1}\|_{2}^{2}-\lambda\left\langle\mathbf{H}_{t}\mathbf{x}_{t+1},\mathbf{x}_{t}-\mathbf{x}_{t+1}\right\rangle,$ (D.6)
where the inequality follows from Assumption B.3. By convexity of $h_{1}(\mathbf{x}):=\frac{1}{2}\mathbf{x}^{\top}\mathbf{H}_{t}\mathbf{x}$, we obtain
@@
We recall that $\mathbf{H}_{t}=\text{diag}(\sqrt{\widehat{\mathbf{v}}_{t}}+\epsilon)$ in Algorithm 2. Combining this with Lemma C.1, we derive
@@
D.3 Proof of Lemma C.1
Proof of Lemma C.1.
@@
$\|\mathbf{v}_{t}\|_{2}$ $=\Big{\|}\sum_{k=1}^{t}(1-\beta_{2,k})\tilde{\mathbf{c}}_{k}^{2}\prod_{j=k+1}^{t}\beta_{2,j}+\prod_{j=1}^{t}\beta_{2,j}\mathbf{v}_{0}\Big{\|}_{2}\leq\sum_{k=1}^{t}(1-\beta_{2,k})\prod_{j=k+1}^{t}\beta_{2,j}=\Big{(}1-\prod_{k=1}^{t}\beta_{2,k}\Big{)}\leq 1,$ (D.7)
@@
$\|\mathbf{m}_{t}\|_{2}$ $=\Big{\|}\sum_{k=1}^{t}(1-\beta_{1,k})\tilde{\mathbf{c}}_{k}\prod_{j=k+1}^{t}\beta_{1,j}+\prod_{j=1}^{t}\beta_{1,j}\mathbf{m}_{0}\Big{\|}_{2}\leq\sum_{k=1}^{t}(1-\beta_{1,k})\prod_{j=k+1}^{t}\beta_{1,j}=\Big{(}1-\prod_{k=1}^{t}\beta_{1,k}\Big{)}\leq 1.$ (D.8)
@@
D.4 Proof of Lemma C.2
Proof of Lemma C.2.
@@
$\mathbf{m}_{t+1}-\nabla F(\mathbf{x}_{t+1})$ $=(1-\beta_{1,t+1})\bigg{(}\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\bigg{)}+\beta_{1,t+1}\big{(}\mathbf{m}_{t}-\nabla F(\mathbf{x}_{t})\big{)}$
@@
$\bm{\varepsilon}_{t+1}=(1-\beta_{1,t+1})\big{(}\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\big{)}+\beta_{1,t+1}\bm{\varepsilon}_{t}+\beta_{1,t+1}\big{(}\gamma_{t+1}\bm{\Delta}_{t}-\mathbb{E}\bm{\Delta}_{t}\big{)},$ (D.9)
where the expectation in the last term is taken over the randomness in $\bm{\xi}_{t+1}$. Taking squared norm over both sides of (D.9) and then take expectation over $\bm{\xi}_{t+1}$, we have
$\mathbb{E}\|\bm{\varepsilon}_{t+1}\|_{2}^{2}$ $=\mathbb{E}\|(1-\beta_{1,t+1})\big{(}\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\big{)}+\beta_{1,t+1}\bm{\varepsilon}_{t}+\beta_{1,t+1}\big{(}\gamma_{t+1}\bm{\Delta}_{t}-\mathbb{E}\bm{\Delta}_{t}\big{)}\|_{2}^{2}$
@@
$+\underbrace{\beta_{1,t+1}^{2}(\gamma_{t+1}-1)^{2}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}}_{\mbox{$II$}}$
$-2(1-\gamma_{t+1})\beta_{1,t+1}\underbrace{\mathbb{E}\bigg{\langle}\bm{\Delta}_{t},(1-\beta_{1,t+1})\big{(}\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\big{)}+\beta_{1,t+1}\bm{\varepsilon}_{t}+\beta_{1,t+1}\big{(}\bm{\Delta}_{t}-\mathbb{E}\bm{\Delta}_{t}\big{)}\bigg{\rangle}}_{\mbox{$III$}}.$ (D.10)
@@
$\leq 2(1-\beta_{1,t+1})^{2}\mathbb{E}\|\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\|_{2}^{2}+2\beta_{1,t+1}^{2}\mathbb{E}\|\bm{\Delta}_{t}-\mathbb{E}\bm{\Delta}_{t}\|_{2}^{2}+\beta_{1,t+1}^{2}\mathbb{E}\|\bm{\varepsilon}_{t}\|_{2}^{2}.$ (D.11)
Minimizing $II-2(1-\gamma_{t+1})\beta_{1,t+1}III$ over $\gamma_{t+1}$ in (D.10), we know that when
$1-\gamma_{t+1}$ $=\frac{\mathbb{E}\bigg{\langle}\bm{\Delta}_{t},(1-\beta_{1,t+1})\big{(}\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\big{)}+\beta_{1,t+1}\bm{\varepsilon}_{t}+\beta_{1,t+1}\big{(}\bm{\Delta}_{t}-\mathbb{E}\bm{\Delta}_{t}\big{)}\bigg{\rangle}}{\beta_{1,t+1}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}},$
$II-2(1-\gamma_{t+1})\beta_{1,t+1}III$ reaches optimality at
$II-2(1-\gamma_{t+1})\beta_{1,t+1}III=-\frac{\Bigg{(}\mathbb{E}\bigg{\langle}\bm{\Delta}_{t},{\color[rgb]{0,0,1}(1-\beta_{1,t+1})\big{(}\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})\big{)}+\beta_{1,t+1}\bm{\varepsilon}_{t}}+\beta_{1,t+1}\big{(}\bm{\Delta}_{t}-\mathbb{E}\bm{\Delta}_{t}\big{)}\bigg{\rangle}\Bigg{)}^{2}}{\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}}.$
Using $G_{t}$ to represent
${\color[rgb]{0,0,1}G_{t+1}:=(1-\beta_{1,t+1}){\color[rgb]{0,0,0}\mathbb{E}\bigg{\langle}\bm{\Delta}_{t},}\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1}){\color[rgb]{0,0,0}\bigg{\rangle}}+\beta_{1,t+1}{\color[rgb]{0,0,0}\mathbb{E}\bigg{\langle}\bm{\Delta}_{t}},\bm{\varepsilon}_{t}{\color[rgb]{0,0,0}\bigg{\rangle}}},$
and defining
$A_{t+1}:=\frac{\Big{(}{\color[rgb]{0,0,1}G_{t+1}}+\beta_{1,t+1}\big{(}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}-\|\mathbb{E}\bm{\Delta}_{t}\|_{2}^{2}\big{)}\Big{)}}{\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}}.$
We have
$II-2(1-\gamma_{t+1})\beta_{1,t+1}III$ $=\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}\Bigg{(}\beta_{1,t+1}(1-\gamma_{t+1})-A_{t+1}\Bigg{)}^{2}-\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}A_{t+1}^{2}$
and when
$\gamma_{t+1}=1-\frac{\Big{(}G_{t+1}+\beta_{1,t+1}\big{(}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}-\|\mathbb{E}\bm{\Delta}_{t}\|_{2}^{2}\big{)}\Big{)}}{\beta_{1,t+1}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}}=\frac{\Big{(}\|\mathbb{E}\bm{\Delta}_{t}\|_{2}^{2}\big{)}-G_{t+1}\Big{)}}{\beta_{1,t+1}\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}}.$ (D.12)
$II-2(1-\gamma_{t+1})\beta_{1,t+1}III$ $=-\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}A_{t+1}^{2}.$
Defining
$M_{t+1}:=\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}A_{t+1}^{2}-\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}\Bigg{(}\beta_{1,t+1}(1-\gamma_{t+1})-A_{t+1}\Bigg{)}^{2}$
we conclude that
@@
$\leq 2(1-\beta_{1,t+1})^{2}\sigma^{2}+2\beta_{1,t+1}^{2}L^{2}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}+\beta_{1,t+1}^{2}\mathbb{E}\|\bm{\varepsilon}_{t}\|_{2}^{2}-M_{t+1}.$
where in the last inequality we utilized the fact that $\mathbb{E}\|\bm{\Delta}_{t}-\mathbb{E}\bm{\Delta}_{t}\|_{2}^{2}\leq\mathbb{E}\|\bm{\Delta}_{t}\|_{2}^{2}$ and the $L$-smoothness of $f$. Rearranging the terms finishes the proof. ∎
D.5 Proof of Lemma C.3 and C.4
Proof of Lemma C.3.
Given the L-smoothness of $F(\mathbf{x})$ in Assuption B.2, we have
@@
$=F(\mathbf{x}_{t})+\left\langle\mathbf{m}_{t},\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\rangle+\left\langle\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t},\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{L}{2}\cdot\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}.$ (D.13)
$\langle\mathbf{m}_{t},\mathbf{x}_{t}-\mathbf{x}_{t+1}\rangle$ $=\big{\langle}\frac{1}{\eta_{t}}\cdot\mathbf{H}_{t}[\mathbf{x}_{t}-\mathbf{x}_{t+1}],\mathbf{x}_{t}-\mathbf{x}_{t+1}\big{\rangle}$
$\geq\frac{\rho}{\eta_{t}}\|\mathbf{x}_{t}-\mathbf{x}_{t+1}\|_{2}^{2}.$ (D.14)
Here the inequality follows from Assumption B.3. ∎
Proof of Lemma C.4.
Bringing (D.14) into (D.13), we obtain
@@
The second inequality follows from applying both the Cauchy-Schwarz and Young’s inequalities. Moreover, the final inequality results from selecting $\eta_{t}$ to satisfy $\eta_{t}<\frac{\rho}{2L}$ This completes our proof.
Bringing (D.5) in Lemma D.2 into (D.13), we have
@@
$-\frac{\rho(1-\eta_{t}\lambda)}{\eta_{t}}\|\mathbf{x}_{t}-\mathbf{x}_{t+1}\|_{2}^{2}-\frac{\lambda}{2}\big{[}\mathbf{x}_{t}^{\top}\mathbf{H}_{t+1}\mathbf{x}_{t}-\mathbf{x}_{t+1}^{\top}\mathbf{H}_{t}\mathbf{x}_{t+1}-\sqrt{2(1-\beta_{2,t})}D^{2}\big{]}.$
@@
$-\frac{\lambda}{2}\big{[}\mathbf{x}_{t}^{\top}\mathbf{H}_{t+1}\mathbf{x}_{t}-\mathbf{x}_{t+1}^{\top}\mathbf{H}_{t}\mathbf{x}_{t+1}-\sqrt{2(1-\beta_{2,t})}D^{2}\big{]}.$
@@
$-\frac{\lambda}{2}\big{[}\mathbf{x}_{t}^{\top}\mathbf{H}_{t+1}\mathbf{x}_{t}-\mathbf{x}_{t+1}^{\top}\mathbf{H}_{t}\mathbf{x}_{t+1}-\sqrt{2(1-\beta_{2,t})}D^{2}\big{]}.$
$=F(\mathbf{x}_{t})+\frac{\eta_{t}}{\rho}\|\nabla F(\mathbf{x}_{t})-\mathbf{m}_{t}\|_{2}^{2}-\frac{\rho}{4\eta_{t}}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|_{2}^{2}-\frac{\lambda}{2}\big{[}\mathbf{x}_{t}^{\top}\mathbf{H}_{t+1}\mathbf{x}_{t}-\mathbf{x}_{t+1}^{\top}\mathbf{H}_{t}\mathbf{x}_{t+1}\big{]}+\frac{\lambda}{2}\sqrt{2(1-\beta_{2,t})}D^{2}.$
@@
D.6 Proof of Lemma C.5
Proof of Lemma C.5.